OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3.

OpenStudy (anonymous):

y=5x-3

OpenStudy (andriod09):

Hey benji! Welcome to Open Study! The first thing you need to do, is use these two formulas and plug in all of the variables. The first one, the change of y, over the change of x, which is \[\frac{\Delta y}{\Delta x}\] and the second one is \[y=mx+b\] So, then you would fill inthe first formula, which is: \[\frac{\Delta y}{\Delta x}= \frac{\Delta y}{\Delta 5x}\] then you would put it into the slop-intercept formula: \y=[mx+b\] like so: \[y=5x-3\] I hope I helped, if you have any further questions about Open Study, how it works, the Code of Conduct, anything at all, feel free to PM a moderator, someone who's name is in full purple, or one with a purple (A) like mine, and we'll get to your question as soon as possible! Have a great day! \[\color{red}{Andy}\color{blue}{|}\color{purple}{(A)Andriod09}\]